OpenStudy (anonymous):

An auditorium earned 25,000 in sold-out concert ticket sales. Front section tickets cost 75$ per seat and back section tickets cost 50$ per seat. The number of front section seats is twice the number of back section seats. How many seats are in the front section?

jimthompson5910 (jim_thompson5910):

first define your variables

jimthompson5910 (jim_thompson5910):

let f = # of front section seats b = # of back section seats

jimthompson5910 (jim_thompson5910):

if the "auditorium earned 25,000 in sold-out concert ticket sales", then we know that 75f + 50b = 25000 and if "The number of front section seats is twice the number of back section seats", then we can say f = 2b ---------------------- 75f + 50b = 25000 ... move back to equation 1 75(2b) + 50b = 25000 ... replace f with 2b (since f = 2b) 150b + 50b = 25000 200b = 25000 b = ???

OpenStudy (anonymous):

But my teacher had given me the answer and said it was 250. and what you did equals to 125.. i think it was f=2b?

jimthompson5910 (jim_thompson5910):

the answer you want is f, not b you found that b = 125 but you have to plug that into f = 2b to find f

OpenStudy (anonymous):

ooooh!

OpenStudy (anonymous):

f=2(125) ?

jimthompson5910 (jim_thompson5910):

yep, so that's why f = 250

OpenStudy (anonymous):

:DD haha thank you!

jimthompson5910 (jim_thompson5910):

you're welcome